CAN ANY ONE HELP TO SOLVE THIS PROBLEM SPECIALLY THE LAST PART. FIND i2?before you come to lab you have to use a similar a
Answer #1

Junction E: i_2-i_1-i_3 =0

Loop BCDE: i_3R_3 - \varepsilon _{2} + i_2R_2=0

Loop ACDF: -i_1R_1+i_3R_3 - \varepsilon _{2} +\varepsilon _{1}=0

Replacing i1 = i2 - i3 in loop ABEF equation

=>\varepsilon _{1} - (i_2 - i_3)R_1 - i_2R_2 = 0

=> \varepsilon _{1} + i_3R_1 - i_2(R_1+R_2) = 0

Solving this equation with loop BCDE simultaneously

=>\frac{\varepsilon _{2}-i_2R_2}{R_3} = \frac{i_2(R_1+R_2)-\varepsilon _{1}}{R_1}

=> i_{2} = \frac{R_1\varepsilon _{2}+\varepsilon _{1}R_{3}}{R_1R_2+R_1R_3+R_2R_3}